Financial statements for Askew Industries for 2021 are shown below (in thousands): 2021 Income Statement Net sales $ 8,800 Coplease complete and show work with formulas
Answer #1
1. Inventory turnover ratio = Cost of goods sold / Average inventory = 6250 / 680 9.19
2. Average days in inventory = Days in a year / Inventory turnover ratio = 365 / 9.19 39.72 days
3. Receivables turnover ratio = Net Sales / Average accounts receivables = 8800 / 480 18.33
4. Average collection period = Days in a year / Receivables turnover ratio = 365 / 18.33 19.91 days
5. Asset turnover ratio = Net sales / Average total assets = 8800 / 3540 2.49
6. Profit margin on sales = Net income / Net sales = 192 / 8800 2.18%
7. Return on assets = Net income / Average total assets = 192 / 3540 5.42%
8. Return on equity = Net income / Average shareholders' equity = 192 / 1385 13.86%
9. Equity multiplier = Average total assets / Average shareholders' equity = 3540 / 1385 2.56
10. Return on shareholders' equity (using the DuPont framework) = Asset turnover ratio * Profit margin on sales * Equity multiplier = 2.49 * 2.18% * 2.56 13.90%
Working notes :
Average inventory = ( Beginning inventory + Ending inventory ) / 2 = ( 580 + 780 ) / 2 680
Average account receivables = ( Beginning account receivable + Ending accounts receivable ) / 2 = ( 380 + 580 ) / 2 480
Average total assets = ( Beginning total assets + Ending total assets ) / 2 = ( 3340 + 3740 ) / 2 3540
Dec. 31
2021 2020
Common stock 580 580
Retained earnings 880 730
Shareholders' equity 1460 1310
Average shareholders' equity = ( Beginning shareholders' equity + Ending shareholders' equity ) / 2 = ( 1310 + 1460 ) / 2 1385